### v3

@@ 6.1.1 Bounding generalization error by Rademacher complexity
$\mathbb{P}\left\{\exists\phi\in\mathcal{F}:\left\|f-\phi\right\|_{L^{2}}^{2}-\left(\mathcal{R}_{S,0}(\phi)-\mathcal{R}_{S,0}\left(f\right)\right)\geq\epsilon\left(\alpha+\beta+\left\|f-\phi\right\|_{L^{2}}^{2}\right)\right\}$
$\leq 14\mathcal{N}\left(\frac{\beta\epsilon}{20B},\mathcal{F},M\right)\exp\left(-\frac{\epsilon^{2}(1-\epsilon)\alpha M}{214(1+\epsilon)B^{4}}\right),$
where
$\mathcal{R}_{S,0}(g)=\frac{1}{M}\sum_{i=1}^{M}|g(\boldsymbol{x}_{i})-y_{i}|^{2}$
for $\{y_{i}\}_{i=1}^{M}\subset[-B,B]$ and $\{\boldsymbol{x}_{i}\}_{i=1}^{M}\subset\Omega$.
Remark 6.17.
@@
In the ordinary Lemma of [16], the covering number of $L^{1}$ is used to bound the inequality, which is stronger than our lemma here. Since $L^{\infty}$ is sufficient to obtain our results and to use Lemma 6.12 directly, we utilize the covering number under the $L^{\infty}$-norm here. Additionally, while their setup involves data with noise, our problem assumes noise-free data. However, the method we use can be easily generalized to accommodate noisy data.
Proposition 6.18.
Let $d,N,L,M\in\mathbb{N}_{+}$, $B,C_{1},C_{2}\in\mathbb{R}_{+}$. For any $f\in W^{n,\infty}([a,b]^{d})$ with $\|f\|_{W^{n,\infty}([a,b]^{d})}\leq B$, we have
$\mathbb{E}\mathcal{R}_{D}(\boldsymbol{\theta}_{S})\leq\frac{5136B^{4}(b-a)^{d}}{M}\left[\log\left(14\mathcal{N}\left(\frac{(b-a)^{d}}{80BM},\widetilde{\Phi},M\right)\right)+1\right]$
(73) $+\frac{5136d^{4}B^{4}(b-a)^{d}}{M}\left[\log\left(14\mathcal{N}\left(\frac{(b-a)^{d}}{80dBM},D^{2}\widetilde{\Phi},M\right)\right)+1\right]+3\inf_{\phi\in\widetilde{\Phi}}\|\phi-f\|_{W^{2,\infty}\left([a,b]^{d}\right)}^{2},$
$\mathbb{E}$ is expected responding to $X$, and $X:=\{\boldsymbol{x}_{1},\ldots,\boldsymbol{x}_{M}\}$ is an independent random variables set uniformly distributed on $[a,b]^{d}$.
Proof 6.19.
Due to the definition, we know that
$\mathcal{R}_{D}(\boldsymbol{\theta}_{D})=\inf_{\phi\in\widetilde{\Phi}}\|\phi-f\|_{H^{2}\left([a,b]^{d}\right)}^{2}.$
For the sample error, due to $\boldsymbol{\theta}_{S}$ and $\boldsymbol{\theta}_{D}$ belong to $\Theta$ almost surely, we have
$\mathbb{E}\left[\mathcal{R}_{D}(\boldsymbol{\theta}_{S})-\mathcal{R}_{S}(\boldsymbol{\theta}_{S})\right]$
$\leq$ $\mathbb{E}\left[\int_{[a,b]^{d}}\left|\nabla f(\boldsymbol{x})-\nabla\phi(\boldsymbol{x};\boldsymbol{\theta}_{S})\right|^{2}\,\mathrm{d}\boldsymbol{x}-\frac{(b-a)^{d}}{M}\sum_{i=1}^{M}\left|\nabla f(\boldsymbol{x}_{i})-\nabla\phi(\boldsymbol{x}_{i};\boldsymbol{\theta}_{S})\right|^{2}\right]$
(74) $+\mathbb{E}\left[\int_{[a,b]^{d}}\left|f(\boldsymbol{x})-\phi(\boldsymbol{x};\boldsymbol{\theta}_{S})\right|^{2}\,\mathrm{d}\boldsymbol{x}-\frac{(b-a)^{d}}{M}\sum_{i=1}^{M}\left|f(\boldsymbol{x}_{i})-\phi(\boldsymbol{x}_{i};\boldsymbol{\theta}_{S})\right|^{2}\right]$
For simplicity, we will only bound the first part; the second part can be bounded in a similar manner.
For the first sample error, we have
$\mathbb{E}\left[\int_{[a,b]^{d}}\left|\Delta f(\boldsymbol{x})-\Delta\phi(\boldsymbol{x};\boldsymbol{\theta}_{S})\right|^{2}\,\mathrm{d}\boldsymbol{x}-\frac{(b-a)^{d}}{M}\sum_{i=1}^{M}\left|\Delta f(\boldsymbol{x}_{i})-\Delta\phi(\boldsymbol{x}_{i};\boldsymbol{\theta}_{S})\right|^{2}\right]$
$=$ $\mathbb{E}\left[\int_{[a,b]^{d}}\left|\Delta f(\boldsymbol{x})-\Delta\phi(\boldsymbol{x};\boldsymbol{\theta}_{S})\right|^{2}\,\mathrm{d}\boldsymbol{x}-\frac{2(b-a)^{d}}{M}\sum_{i=1}^{M}\left|\Delta f(\boldsymbol{x}_{i})-\Delta\phi(\boldsymbol{x}_{i};\boldsymbol{\theta}_{S})\right|^{2}\right]+\mathbb{E}\frac{(b-a)^{d}}{M}\sum_{i=1}^{M}\left|\Delta f(\boldsymbol{x}_{i})-\Delta\phi(\boldsymbol{x}_{i};\boldsymbol{\theta}_{S})\right|^{2}$
(75) $\leq$ $\mathbb{E}\left[\int_{[a,b]^{d}}\left|\Delta f(\boldsymbol{x})-\Delta\phi(\boldsymbol{x};\boldsymbol{\theta}_{S})\right|^{2}\,\mathrm{d}\boldsymbol{x}-\frac{2(b-a)^{d}}{M}\sum_{i=1}^{M}\left|\Delta f(\boldsymbol{x}_{i})-\Delta\phi(\boldsymbol{x}_{i};\boldsymbol{\theta}_{S})\right|^{2}\right]+\inf_{\phi\in\widetilde{\Phi}}\|\phi-f\|_{W^{2,\infty}\left([a,b]^{d}\right)}^{2}$
where the last inequality is due to the definition of $\boldsymbol{\theta}_{S}$.
@@
(76) $\mathbb{A}_{1}=\int_{[a,b]^{d}}\left|\Delta f(\boldsymbol{x})-\Delta\phi(\boldsymbol{x};\boldsymbol{\theta}_{S})\right|^{2}\,\mathrm{d}\boldsymbol{x}-\frac{2(b-a)^{d}}{M}\sum_{i=1}^{M}\left|\Delta f(\boldsymbol{x}_{i})-\Delta\phi(\boldsymbol{x}_{i};\boldsymbol{\theta}_{S})\right|^{2}.$
Note
$\mathbb{P}(\mathbb{A}_{1}\geq\epsilon)$ $=\mathbb{P}\left(2\|\Delta f-\Delta\phi\|^{2}_{L^{2}([a,b]^{d})}-\frac{2(b-a)^{d}}{M}\sum_{i=1}^{M}\left|\Delta f(\boldsymbol{x}_{i})-\Delta\phi(\boldsymbol{x}_{i};\boldsymbol{\theta}_{S})\right|^{2}\geq\epsilon+\|\Delta f-\Delta\phi\|^{2}_{L^{2}([a,b]^{d})}\right)$
(77) $\leq 14\mathcal{N}\left(\frac{\epsilon}{80dB},D^{2}\widetilde{\Phi},M\right)\exp\left(-\frac{\epsilon M(b-a)^{d}}{5136d^{4}B^{4}}\right),$
where the last inequality is due to Lemma 6.16.
Therefore, we have that
$\mathbb{E}\mathbb{A}_{1}$ $\leq\int_{0}^{\infty}\mathbb{P}(\mathbb{A}_{1}\geq y)\,\mathrm{d}y\leq\epsilon+\int_{\epsilon}^{\infty}\mathbb{P}(\mathbb{A}_{1}\geq y)\,\mathrm{d}y$
(78) $\leq\epsilon+\int_{\epsilon}^{\infty}14\mathcal{N}\left(\frac{\epsilon}{80dB},D^{2}\widetilde{\Phi},M\right)\exp\left(-\frac{yM}{5136d^{4}B^{4}(b-a)^{d}}\right)\,\mathrm{d}y.$
By the direct calculation, we have
$\int_{\epsilon}^{\infty}14\mathcal{N}\left(\frac{\epsilon}{80dB},D^{2}\widetilde{\Phi},M\right)\exp\left(-\frac{yM}{5136d^{4}B^{4}(b-a)^{d}}\right)\,\mathrm{d}y$
(79) $\leq$ $14\mathcal{N}\left(\frac{\epsilon}{80dB},D^{2}\widetilde{\Phi},M\right)\frac{5136d^{4}B^{4}(b-a)^{d}}{M}\exp\left(-\frac{\epsilon M}{5136d^{4}B^{4}(b-a)^{d}}\right).$
Set
$\epsilon=\frac{5136d^{4}B^{4}(b-a)^{d}}{M}\log\left(14\mathcal{N}\left(\frac{(b-a)^{d}}{80dBM},D^{2}\widetilde{\Phi},M\right)\right)\geq\frac{(b-a)^{d}}{M}$
@@
$\mathbb{E}\mathbb{A}_{1}\leq\frac{5136B^{4}d^{4}(b-a)^{d}}{M}\left[\log\left(14\mathcal{N}\left(\frac{(b-a)^{d}}{80BM},D^{2}\widetilde{\Phi},M\right)\right)+1\right]$
@@
$\mathbb{E}\left[\mathcal{R}_{D}(\boldsymbol{\theta}_{S})-\mathcal{R}_{S}(\boldsymbol{\theta}_{S})\right]\leq\frac{5136B^{4}(b-a)^{d}}{M}\left[\log\left(14\mathcal{N}\left(\frac{(b-a)^{d}}{80BM},\widetilde{\Phi},M\right)\right)+1\right]$
(80) $+\frac{5136d^{4}B^{4}(b-a)^{d}}{M}\left[\log\left(14\mathcal{N}\left(\frac{(b-a)^{d}}{80dBM},D^{2}\widetilde{\Phi},M\right)\right)+1\right]+2\inf_{\phi\in\widetilde{\Phi}}\|\phi-f\|_{W^{2,\infty}\left([a,b]^{d}\right)}^{2}.$
Proof 6.20 (Proof of Theorem 6.2).
The proof can be obtained directly by combining Proposition 6.18, Lemma 6.12, and Theorem 6.13.

### v4

@@ 6 Generalization Analysis in Sobolev Spaces
$\mathbb{P}\left\{\exists\phi\in\mathcal{F}:\left\|f-\phi\right\|_{L^{2}}^{2}-\left(\mathcal{E}_{\mathcal{S}}(\phi)-\mathcal{E}_{\mathcal{S}}\left(g\right)\right)\geq\epsilon\left(\alpha+\beta+\left\|g-\phi\right\|_{L^{2}}^{2}\right)\right\}$
[…]
@@
In the ordinary Lemma of [23], the covering number of $l^{1}$ is used to bound the inequality, which is stronger than our lemma here. Since $l^{\infty}$ is sufficient to obtain our results and to use Lemma 17 directly, we utilize the covering number under the $l^{\infty}$-norm here.
With the preparations, we are ready to bound the generalization error by combining the approximation error with a covering-number estimate, as stated in the following proposition:
Proposition 16.
Let $M\in\mathbb{N}$, and assume that $\|\mathcal{L}f_{\rho}\|_{L^{\infty}(\Omega)}\leq B$ and $\mathcal{Y}\in[-L,L]$ are almost surely for some $B\geq 1$. Then we have that
$\mathbb{E}\|\mathcal{L}f_{\mathcal{S},\mathcal{F}_{N,L,D_{1},D_{2},D_{3},B}}-\mathcal{L}f_{\rho}\|_{L^{2}(\Omega)}^{2}\leq 2\inf_{f\in\mathcal{F}_{N,L,D_{1},D_{2},D_{3},B}}\left\|\mathcal{L}f-\mathcal{L}f_{\rho}\right\|_{L^{2}(\Omega)}^{2}$
$+$ $\frac{5136B^{4}}{M}\left\{\log\left(14\ \mathcal{N}\left(\frac{1}{80BM},\mathcal{L}\mathcal{F}_{N,L,D_{1},D_{2},D_{3},B},M\right)\right)+1\right\}.$
Proof.
To simplicity notations, we denote $\mathcal{F}_{N,L,D_{1},D_{2},D_{3},B}$ as $\mathcal{F}_{*}$ in this proof. First, we divide $\|\mathcal{L}f_{\mathcal{S},\mathcal{F}_{*}}-\mathcal{L}f_{\rho}\|_{L^{2}(\Omega)}^{2}$ into two parts
$\|\mathcal{L}f_{\mathcal{S},\mathcal{F}_{*}}-\mathcal{L}f_{\rho}\|_{L^{2}(\Omega)}^{2}=2(\mathcal{E}_{\mathcal{S}}(f_{\mathcal{S},\mathcal{F}_{*}})-\mathcal{E}_{\mathcal{S}}(f_{\rho}))$
$+$ $\underbrace{\|\mathcal{L}f_{\mathcal{S},\mathcal{F}_{*}}-\mathcal{L}f_{\rho}\|_{L^{2}(\Omega)}^{2}-2(\mathcal{E}_{\mathcal{S}}(f_{\mathcal{S},\mathcal{F}_{*}})-\mathcal{E}_{\mathcal{S}}(f_{\rho}))}_{\mathbf{A}_{0}}.$
For the first part, we have that
$2\mathbb{E}(\mathcal{E}_{\mathcal{S}}(f_{\mathcal{S},\mathcal{F}_{*}})-\mathcal{E}_{\mathcal{S}}(f_{\rho}))\leq 2\mathbb{E}\inf_{f\in\mathcal{F}_{*}}(\mathcal{E}_{\mathcal{S}}(f)-\mathcal{E}_{\mathcal{S}}(f_{\rho}))\leq 2\inf_{f\in\mathcal{F}_{*}}\mathbb{E}(\mathcal{E}_{\mathcal{S}}(f)-\mathcal{E}_{\mathcal{S}}(f_{\rho}))$
$=$ $2\inf_{f\in\mathcal{F}_{*}}\mathbb{E}\left(\frac{1}{M}\sum_{j=1}^{M}|\mathcal{L}f(\bm{x}_{i})-\mathcal{L}f_{\rho}(\bm{x}_{i})|^{2}\right)\leq 2\inf_{f\in\mathcal{F}_{*}}\left\|\mathcal{L}f-\mathcal{L}f_{\rho}\right\|_{L^{2}(\Omega)}^{2},$
where the first equality is use to
$\mathbb{E}_{\bm{x}_{i}}(\mathcal{L}f(\bm{x}_{i})-\mathcal{L}f_{\rho}(\bm{x}_{i}))\mathbb{E}_{y_{i}}[(\mathcal{L}f_{\rho}(\bm{x}_{i})-y_{i})|\bm{x}_{i}]=\mathbb{E}_{\bm{x}_{i}}(\mathcal{L}f(\bm{x}_{i})-\mathcal{L}f_{\rho}(\bm{x}_{i}))\cdot 0=0,$
and last inequality is due to Assumption 1. As for $\mathbf{A}_{0}$, we have
$\quad\mathbb{P}(\mathbf{A}_{0}\geq\epsilon)$
$=\mathbb{P}\left(2\|\mathcal{L}f_{\mathcal{S},\mathcal{F}_{*}}-\mathcal{L}f_{\rho}\|_{L^{2}(\Omega)}^{2}-2(\mathcal{E}_{\mathcal{S}}(f_{\mathcal{S},\mathcal{F}_{*}})-\mathcal{E}_{\mathcal{S}}(f_{\rho}))\geq\epsilon+\|\mathcal{L}f_{\mathcal{S},\mathcal{F}_{*}}-\mathcal{L}f_{\rho}\|_{L^{2}(\Omega)}^{2}\right)$
$=\mathbb{P}\left(\|\mathcal{L}f_{\mathcal{S},\mathcal{F}_{*}}-\mathcal{L}f_{\rho}\|_{L^{2}(\Omega)}^{2}-(\mathcal{E}_{\mathcal{S}}(f_{\mathcal{S},\mathcal{F}_{*}})-\mathcal{E}_{\mathcal{S}}(f_{\rho}))\geq\frac{1}{2}\left(\epsilon+\|\mathcal{L}f_{\mathcal{S},\mathcal{F}_{*}}-\mathcal{L}f_{\rho}\|_{L^{2}(\Omega)}^{2}\right)\right)$
$\leq 14\mathcal{N}\left(\frac{\epsilon}{80B},\mathcal{L}\mathcal{F}_{*},M\right)\exp\left(-\frac{\epsilon M}{5136B^{4}}\right),$
where the last inequality is due to Lemma 15 for choosing $\alpha=\beta=\frac{1}{2}\epsilon$ and $\delta=\frac{1}{2}$.
Therefore, we have that
$\mathbb{E}\mathbf{A}_{0}$ $\leq\int_{0}^{\infty}\mathbb{P}(\mathbf{A}_{0}\geq t)\,\mathrm{d}t\leq\epsilon+\int_{\epsilon}^{\infty}\mathbb{P}(\mathbf{A}_{0}\geq t)\,\mathrm{d}y$
[…]
By the direct calculation, we have
$\int_{\epsilon}^{\infty}14\mathcal{N}\left(\frac{\epsilon}{80B},\mathcal{L}\mathcal{F}_{*},M\right)\exp\left(-\frac{tM}{5136B^{4}}\right)\,\mathrm{d}t$
$\leq$ $14\mathcal{N}\left(\frac{\epsilon}{80B},\mathcal{L}\mathcal{F}_{*},M\right)\frac{5136B^{4}}{M}\exp\left(-\frac{\epsilon M}{5136B^{4}}\right).$
@@
$\epsilon=\frac{5136B^{4}}{M}\log\left(14\mathcal{N}\left(\frac{1}{80BM},\mathcal{L}\mathcal{F}_{*},M\right)\right)\geq\frac{1}{M}$
@@
$\mathbb{E}\mathbf{A}_{0}\leq\frac{5136B^{4}}{M}\left[\log\left(14\mathcal{N}\left(\frac{1}{80BM},\mathcal{L}\mathcal{F}_{*},M\right)\right)+1\right]$
@@
$\mathbb{E}\|\mathcal{L}f_{\mathcal{S},\mathcal{F}_{*}}-\mathcal{L}f_{\rho}\|_{L^{2}(\Omega)}^{2}$
$\leq$ $\frac{5136B^{4}}{M}\left[\log\left(14\mathcal{N}\left(\frac{1}{80BM},\mathcal{L}\mathcal{F}_{*},M\right)\right)+1\right]+2\inf_{f\in\mathcal{F}_{*}}\left\|\mathcal{L}f-\mathcal{L}f_{\rho}\right\|_{L^{2}(\Omega)}^{2}.$
We bound the covering numbers via the pseudo-dimension, defined below.
Definition 9 (pseudo-dimension [47]).
Let $\mathcal{F}$ be a class of functions from $\mathcal{X}$ to $\mathbb{R}$. The pseudo-dimension of $\mathcal{F}$, denoted by $\operatorname{Pdim}(\mathcal{F})$, is the largest integer $m$ for which there exists $(x_{1},x_{2},\ldots,x_{m},y_{1},y_{2},\ldots,y_{m})\in\mathcal{X}^{m}\times\mathbb{R}^{m}$ such that for any $(b_{1},\ldots,b_{m})\in\{0,1\}^{m}$ there is $f\in\mathcal{F}$ such that $\forall i:f\left(x_{i}\right)>y_{i}\Longleftrightarrow b_{i}=1.$
Lemma 17 ([2], Theorem 12.2).
Let $\mathcal{F}$ be a class of functions from $\mathcal{X}$ to [B,B][-B,B]. For any $\varepsilon>0$, we have
$\mathcal{N}(\varepsilon,\mathcal{F},n)\leq\left(\frac{2\mathrm{e}nB}{\varepsilon\operatorname{Pdim}(\mathcal{F})}\right)^{\operatorname{Pdim}(\mathcal{F})}$
for $n\geq\operatorname{Pdim}(\mathcal{F})$.
Proposition 18.
For any $N,L,d,D_{i},B\in\mathbb{N}_{+}$, there exists a constant $\bar{C}$ independent of $N,L$ such that
$\operatorname{Pdim}(\mathcal{L}\mathcal{F}_{N,L,D_{1},D_{2},D_{3},B})\leq\bar{C}N^{2}L^{2}\log_{2}L\log_{2}N,$
where $\mathcal{L}\mathcal{F}_{N,L,D_{1},D_{2},D_{3},B}$ is defined in (6.1) and $\mathcal{L}$ satisfies Assumption 1.
Proof.
Denote
$\Phi_{\mathcal{N}}:=\{\eta(\bm{x},y):\eta(\bm{x},y)=\psi(\bm{x})-y,\;\psi\in\mathcal{L}\mathcal{F}_{N,L,D_{1},D_{2},D_{3},B},\;(\bm{x},y)\in\mathbb{R}^{d+1}\}.$
By the definitions of VC-dimension and pseudo-dimension, we have
$\operatorname{Pdim}(\mathcal{L}\mathcal{F}_{N,L,D_{1},D_{2},D_{3},B})\;\leq\;\operatorname{VCdim}(\Phi_{\mathcal{N}}).$ (6.2)
For $\operatorname{VCdim}(\Phi_{\mathcal{N}})$, it can be bounded by $\mathcal{O}(N^{2}L^{2}\log_{2}L\log_{2}N)$. The proof is analogous to that for the estimate of $\operatorname{VCdim}(D^{2}\Phi)$ given in Theorem 10. The form of $\mathcal{L}$ does not affect the proof since $\mathcal{L}$ is a linear operator; for any fixed input $\bm{x}$, the output remains a piecewise polynomial with respect to the parameter space. Therefore, the argument used for $\operatorname{VCdim}(D^{2}\Phi)$ in Theorem 10 still applies here. ∎
Proof of Theorem 14.
Based on Lemma 17, Propositions 16 and 18, set $M_{0}=\operatorname{Pdim}(\mathcal{L}\mathcal{F}_{N,L,D_{1},D_{2},D_{3},B})$, we have that for any $M\geq M_{0}$,
$\mathbb{E}\|\mathcal{L}f_{\mathcal{S},\mathcal{F}_{N,L,D_{1},D_{2},D_{3},B}}-\mathcal{L}f_{\rho}\|_{L^{2}(\Omega)}^{2}$
$\leq$ $\frac{5136B^{4}}{M}\left[\log\left(14\mathcal{N}\left(\frac{1}{80BM},\mathcal{L}\mathcal{F}_{N,L,D_{1},D_{2},D_{3},B},M\right)\right)+1\right]+C_{*}(NL)^{-\frac{4(n-2)}{d}}$
$\leq$ $\frac{5136B^{4}}{M}\left[\log\left(\mathcal{N}\left(\frac{1}{80BM},\mathcal{L}\mathcal{F}_{N,L,D_{1},D_{2},D_{3},B},M\right)\right)+4\right]+C_{*}(NL)^{-\frac{4(n-2)}{d}}$
$\leq$ $\frac{5136B^{4}}{M}\left[\operatorname{Pdim}(\mathcal{L}\mathcal{F}_{N,L,D_{1},D_{2},D_{3},B})\log\left(160\mathrm{e}M^{2}B^{2}\right)+4\right]+C_{*}(NL)^{-\frac{4(n-2)}{d}}$
$\leq$ $\frac{5136B^{4}}{M}\left[\bar{C}N^{2}L^{2}\log_{2}N\log_{2}L\log\left(160\mathrm{e}M^{2}B^{2}\right)+4\right]+C_{*}(NL)^{-\frac{4(n-2)}{d}}$
$=$ $\mathcal{O}\left(\frac{N^{2}L^{2}\log N\log L\log M}{M}+(NL)^{-\frac{4(n-2)}{d}}\right).$
We consider
$E_{1}\;=\;\frac{N^{2}L^{2}\,(\log N)(\log L)\,\\ logM}{M},\qquad E_{2}\;=\;(NL)^{-\frac{4(n-2)}{d}}.$
Let $P:=NL$ and assume $N=\mathcal{O}(L)$, so that $(\log N)(\log L)=\mathcal{O}\left((\log P)^{2}\right)$. Balancing the two terms, $E_{1}=E_{2}$, yields
$P^{\,2+\alpha}(\log P)^{2}\;=\;\mathcal{O}\left(\frac{M}{\log M}\right),\qquad\alpha:=\frac{4(n-2)}{d}.$
Solving for the optimal $P$ gives
$P^{\star}=\mathcal{O}\left(\;M^{\frac{1}{2+\alpha}}\,(\log M)^{-\frac{3}{2+\alpha}}\right)\quad\text{(ignoring lower-order $\log\log M$ factors)}.$
At the balance point, the minimal error satisfies
$M^{-\frac{2(n-2)}{\,d+2(n-2)\,}}\,(\log M)^{\frac{6(n-2)}{\,d+2(n-2)\,}}.$